Its evaluate the integrals using these values... Can someone show me how to do this step by step? Please and Thank you!
ļ rat = 8 | dt=4 {&a=64, $(2x +5x+10)dx
Answer #1

Given that ſp ds = 64.med = Band Jis = 4 Consider ( 2x +53 +10) dr = 2{xdx +s+10dr= 2x64+5*8*10*4 = 128+40 +40 – 208)
Answer #2

SOLUTION :


 4

∫ (2 x^3 + 5 x + 10) dx

0


       4                    4                  4

= 2 ∫. x^3 dx  + 5 ∫. X dx  + 10 ∫  dx

      0                    0                   0


= 2 * 64  + 5 * 8 + 10 * 4


= 128 + 40 + 40


= 208 (ANSWER)

answered by: Tulsiram Garg